### v2

# PSHuman: Photorealistic Single-image 3D Human Reconstruction using Cross-Scale Multiview Diffusion and Explicit Remeshing

Abstract
Photorealistic 3D human modeling is essential for various applications and has seen tremendous progress. However, existing methods for monocular full-body reconstruction, typically relying on front and/or predicted back view, still struggle with satisfactory performance due to the ill-posed nature of the problem and sophisticated self-occlusions. In this paper, we propose PSHuman, a novel framework that explicitly reconstructs human meshes utilizing priors from the multiview diffusion model. It is found that directly applying multiview diffusion on single-view human images leads to severe geometric distortions, especially on generated faces. To address it, we propose a cross-scale diffusion that models the joint probability distribution of global full-body shape and local facial characteristics, enabling identity-preserved novel-view generation without geometric distortion. Moreover, to enhance cross-view body shape consistency of varied human poses, we condition the generative model on parametric models (SMPL-X), which provide body priors and prevent unnatural views inconsistent with human anatomy. Leveraging the generated multiview normal and color images, we present SMPLX-initialized explicit human carving to recover realistic textured human meshes efficiently. Extensive experiments on CAPE and THuman2.1 demonstrate PSHuman’s superiority in geometry details, texture fidelity, and generalization capability.
1 Introduction
Photorealistic 3D reconstruction of clothed humans is a promising and widely investigated research domain with significant applications across several industries, including gaming, movies, fashion, and AR/VR [26, 29]. Traditional methods, which perform multiview stereo and non-rigid registration using multi-camera setups or incorporate additional depth signals, have achieved accurate modeling. However, reconstruction from an in-the-wild RGB image remains an open problem due to sophisticated body poses and complex clothing topology.
      (a)                  (b)
Figure 1: Each triplet contains input (left) and reconstructions of w/o (middle) and w/ (right) SMPL-X condition. Compared with naive diffusion, SMPL-X prior guides handling self-occlusion and improving consistency.
Early studies [35, 36, 42] utilize implicit functions [27, 31] to recover textured human mesh from a single color or normal image. Despite improvements in monocular ambiguity and postural intricacy, this regression paradigm still falls short in detail loss and novel view artifacts. Recent efforts [51, 13] incorporate generative information, such as predicting a back view, to mitigate these issues. On the one hand, the reconstruction pipelines still follow implicit functions, which exhibit limitations in capturing high-fidelity geometry and texture details. On the other hand, the introduction of the back view fails to provide enough stereo information to mitigate spatial ambiguity.
In this study, we aim to tackle these existing challenges by introducing a multiview diffusion model and a normal-guided explicit human reconstruction framework. Different from the front and/or back views by existing methods [43, 13], we explore the direct multiple views generation for robust human modeling. As depicted in Fig. 2, PSHuman takes a full-body human image as input, followed by a designed multiview diffusion model and an SMPLX-initialized mesh carving module, outputting a textured 3D human mesh.
Specifically, we fine-tune a pre-trained text-to-image diffusion model (such as Stable Diffusion [34]) to generate multiview color and normal maps conditioned on the input reference. Despite impressive generative performance, this base framework faces two major challenges: 1) Unnatural body structures, where diffusion models struggle to generate reasonable novel views of posed human, often resulting in disproportionate body proportions or missing body parts. As shown in Fig. 1, this issue arises from the severe self-occlusion in the posed human image and lack of body prior for generative models. To address this, we propose a SMPL-X conditioned diffusion model, which concatenates renderings of estimated SMPL-X with the input image to provide pose guidance for novel-view generation. This approach constrains the diffusion model to generate consistent views that adhere to human anatomy, even when fine-tuning with as few as $3,000$ human scans. 2) Face distortion, where pre-trained diffusion models often produce distorted and unnatural face details, especially for full-body human input. This problem is attributed to the small size of the face in full-body images, which provides limited information for detailed normal prediction after VAE encoding. To accurately recover face geometry, we propose a body-face cross-scale diffusion framework that simultaneously generates multiview full-body images and local face ones. We also employ a simple yet efficient noise blending layer to enhance face details in global image, guaranteeing both cross-scale and cross-view consistency. Consequently, PSHuman generates high-quality and detailed novel-view human images and corresponding normal maps.
To fully leverage the generated multiview images, we present a SMPLX-initialized explicit human carving module for fast and high-fidelity textured human mesh modeling. Unlike implicit functions that use Multilayer Perceptrons (MLPs) to map normal features to an implicit surface, or BiNI [3] that utilizes variational normal integration to recover 2.5D surfaces, we directly reconstruct the 3D mesh supervised by generated multiview normal maps. In practice, we initialize the human model with predicted SMPL-X, and deform and remesh it with differentiable rasterization [30]. As shown in Fig. LABEL:fig_teaser, PSHuman can preserve fine-grained details, such as facial features and fabric wrinkles, and generate natural and harmonious novel views. For texturing on the generated meshes, we first fuse multiview color images using differentiable rendering to mitigate generative inconsistencies, then project them onto the reconstructed 3D mesh.
The entire reconstruction process takes as short as one minute. It is noted that recent SDS-based methods [15, 14] also achieve state-of-the-art performance in geometry details and appearance fidelity. However, they can only handle simple poses and suffer from time-consuming optimization (e.g., TeCH [15] takes approximately six hours). Conversely, PSHuman achieves a balance between precision, efficiency, and pose robustness.
In summary, our key contributions include:
  • We introduce PSHuman, a novel diffusion-based explicit method for detailed and realistic 3D human modeling from a single image.
  • We present a body-face cross-scale diffusion and a SMPL-X conditioned multiview diffusion for high-quality full-body human image generation with high-fidelity face details.
  • We design a SMPLX-initialized explicit human carving module to fast recover textured human mesh based on generated multiview cross-domain images, achieving SOTA performance on THuman2.1 and CAPE datasets.
2 Related Works
Implicit Human Reconstruction. Implicit functions have gained significant traction in human reconstruction [4, 8, 44] due to their flexibility in handling complex topology and diverse clothing styles. Pioneering works such as PIFu [35] introduce pixel-aligned implicit functions, mapping 2D image features to 3D implicit surface for continuous modeling. Building upon this, subsequent research incorporates parametric models (e.g., SMPL) to enhance anatomical plausibility and robustness in challenging in-the-wild poses [10, 42, 54, 50] or for animation-ready modeling [16, 11]. Other efforts enhance geometric details and dynamic stability by introducing normal [36], depth clues  [47, 52], or decoupling albedo [2] from natural inputs. However, these methods struggle with unseen areas due to limited observed information. More recent approaches [51, 13] incorporate predicted side-view images to enhance visualization but still face challenges in balancing quality, efficiency, and robustness.
Explicit Human Reconstruction. Early research focuses on explicit representation for human reconstruction. Voxel-based methods [39, 53] utilize 3D UNet to predict volumetric confidence occupied by the human body, which demands high memory and often results in compromised spatial resolution, hindering the capture of fine details crucial for realistic representation. As a more efficient alternative, visual hulls [28] approximate 3D shapes by incorporating silhouettes and 3D joints. Another strategy involves using depth  [6, 37, 9] or normal [1, 43] information to explicitly infer the 3D human body, balancing detail preservation with computational efficiency. Among these, ECON utilizes normal integration and shape completion, achieving extreme robustness for challenging poses and loose clothing. The major limitations lie in sub-optimal geometry and supporting appearance. To address this, we propose to simultaneously recover geometry and appearance with differentiable rasterization under the supervision of multiview normal and color maps predicted by the diffusion model.
Diffusion-based Human Reconstruction. Most recently, Score Distillation Sampling (SDS) [32] based human generation methods [22, 15] have achieved SOTA performance. However, these approaches often require time-consuming optimization. Drawing inspiration from the advancement of multiview diffusion based 3D generation [23, 24, 21, 40, 38], our work reduces the inference time by directly generating multiple human views for human reconstruction. We further augment human generation capabilities through the introduction of a novel SMPL-X-conditioned cross-scale attention framework. Most related to our work, Chupa [19] also reconstructs with multiview normals. However, it still depends on optimization-based refinement and does not support image condition and texture modeling.
3 Method
Overview. Given a single color image, PSHuman recovers a textured human mesh by two primary stages: 1) a body-face cross-scale multiview diffusion conditioned on SMPL-X, which generates multiview full-body cross-domain (color and normal) images and local facial ones (Sec. 3.1), 2) an SMPLX-initialized explicit human carving module for modeling 3D textured meshes (Sec. 3.2). Different from previous works utilizing front and/or back views, we follow [21, 24] to directly generate six views (front, front left, left, back, right, and front right) for explicit reconstruction, which strike the best balance between computational cost and effectiveness. Since we generate normal maps and images, we use $x$ and $z$ as the raw data and latent for both modalities.
Figure 2: (a) Overall pipeline. Given a single full-body human image, PSHuman recovers the texture human mesh by two stages: 1) Body-face enhanced and SMPL-X conditioned multiview generation. The input image and predicted SMPL-X are fed into a multiview image diffusion model to generate six views of global full-body images and front local face images. 2) SMPLX-initialized explicit human carving. Utilizing generated normal and color maps to deform and remesh the SMPL-X with differentiable rasterization. (b) Illustration of joint denoising diffusion block.
3.1 Body-face Multiview Diffusion
3.1.1 Body-face Diffusion
Motivation. Simply adopting the multiview diffusion [21, 24] for 3D human reconstruction leads to distorted faces and altered facial identities. Because the face only occupies a small region with a low resolution in the image and cannot be accurately generated by the multiview diffusion model. Since humans are very sensitive to slight changes in faces, such generation inaccuracy of faces leads to obvious distortion and identity changes. This motivates us to separately apply another multiview diffusion model to generate the face at a high resolution with more accuracy.
Forward and reverse processes. We define our data distribution $p(x)$ as the joint distribution of the human face $x^{F}$ and the human body $x^{B}$ by
$p(\mathbf{x})=p(x^{B},x^{F})=p(x^{B}|x^{F})p(x^{F}).$ (1)
Then, we follow the DDPM model to define our forward and reverse diffusion process by
$q(x_{t}|x_{t-1})=q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})q(x^{F}_{t}|x^{F}_{t-1}),$ (2)
$p(x_{t-1}|x_{t})=p(x^{B}_{t-1}|x^{B}_{t},x^{F}_{t-1})p(x^{F}_{t-1}|x^{F}_{t}),$ (3)
where $q$ defines the forward process to add noise to the original data and $p$ defines the reverse process to generate data by denoising. For the forward process, we omit the condition on the $x^{F}_{t-1}$ and add noises to the face and body images separately by the approximated forward process
$q(x_{t}|x_{t-1})\approx q(x^{B}_{t}|x^{B}_{t-1})q(x^{F}_{t}|x^{F}_{t-1}).$ (4)
Although explicitly defining forward process for $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$ is feasible for the vanilla diffusion model, it is difficult for the latent diffusion model. We explain this difficulty and the feasibility of this approximation in supplementary material. For the reverse process $p(x_{t-1}|x_{t})$, the face diffusion is just a vanilla diffusion model $p(x^{F}_{t-1}|p^{F}_{t})$ while the body diffusion model will additionally use the face denoising results as conditions by $p(x^{B}_{t-1}|p^{B}_{t},p^{F}_{t-1})$, as shown in Fig. 2(b), which is implemented by the following joint denoising scheme.
Joint denoising. We utilize a simple but efficient noise blending layer to jointly denoise in body-face diffusion. Specifically, in each self-attention block of UNet, we extract the latent vector of the face branch, resize it with scale $s$, and add it to the face region of the global branch with a weight $w$. Specifically, let us take one of the hidden layers as an example. We denote $h^{B_{n}}_{t}$ and $h^{F}_{t}$ as hidden vectors of the $n$-th body view and face view at the same attention layer  [^1] Here, we omit the layer subscript for simplicity. and timestep $t$, the blending operation can be written as
$h^{B_{n}}_{t}$ =$=\begin{cases}h^{B_{1}}_{t}+w\cdot\mathop{RP}(h^{F}_{t},s),&n=1\cr h^{B_{n}}_{t},&n=2,3,\dots,N\end{cases}$ (5)
where the $\mathop{RP}$ is the resize and padding function, and $w$ is a binary mask of the face region, which is obtained with a face detector or a straightforward cropping strategy. The resulting latent vector can be represented by $z^{B_{n}}_{t}$ and $z^{F}_{t}$. We jointly optimize the body and face distribution with the following loss,
$\ell=$ $~{}~{}~{}~{}\mathbb{E}_{t,{z}^{F}_{0},\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(z^{F}_{t},t)\|_{2}\right]$
$\ \ +\mathbb{E}_{t,{z}^{B}_{0},{z}^{F}_{0},n,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}^{(n)}-\mathbf{\epsilon}^{(n)}_{\theta}(z^{B}_{t},z^{F}_{t},t)\|_{2}\right],$ (6)
where $\theta$ is shared weights between face and body views. The noise blending allows the face information to be transferred to novel body views with cross-view attention, improving the overall consistency of generated human images.
3.1.2 SMPL-X Guided Multiview Diffusion
Our multiview diffusion model excels in generating plausible novel views for simple posed images, producing natural human geometry. However, it faces challenges with in-the-wild images that often feature self-occlusions. These occlusions can lead to “hallucinations” that violate human structural integrity or exhibit inconsistent limb poses. For example, Fig. 1 illustrates two common issues: (a) the model generating upright side views for a bending posture, and (b) inconsistencies in arm regions of side views due to self-occlusion, resulting in failed reconstruction.
To mitigate these impediments, we propose incorporating additional pose guidance into the diffusion process. Our method first estimates the SMPL-X of the input image and renders them from six target viewpoints. We then utilize a pre-trained Variational Autoencoder (VAE) encoder to convert SMPL-X renderings and reference images into latent vectors, which are concatenated with noise samples to serve as input of the denoising UNet. The introduction of these conditional signals constrains the multiview distribution, leading to more accurate and consistent human image generation. This approach significantly enhances the model’s generalization capability on complex human poses with self-occlusion.
3.2 SMPLX-initialized Explicit Human Carving
Figure 3: Illustration of our explicit human carving module.
Following the generation of multiview color and normal images, we elaborate on our SMPLX-initialized human carving module (Fig. 3) to obtain the textured 3D mesh.
Numerous methodologies have been developed to leverage normal cues for human reconstruction. However, a significant proportion of them employ implicit functions (e.g. MLP) to map the normal feature as implicit surfaces. This process, while effective in certain scenarios, often results in a lack of fine geometric details. Even with BiNI used in ECON, the overall geometry still exhibits a notable degradation. Taking advantage of the multiview consistent normal maps, we opt to fuse it directly with the explicit triangle mesh. Our reconstruction module consists of three main stages: SMPL-X initialization, differentiable remeshing, and appearance fusion.
SMPL-X initialization. The process commences with human mesh initialization, utilizing the aforementioned SMPL-X estimation, which provides a strong body prior, effectively mitigating unnecessary face pruning and densification during subsequent geometry optimization. However, it is noteworthy that the generated multiple views may exhibit slight misalignment with the SMPL model due to normalization and recentering procedures. Drawing inspiration from ICON, we optimize SMPL-X’s translation $\tilde{t}$, shape $\tilde{\beta}$, and pose $\tilde{\theta}$, parameters to minimize:
$\tilde{t},\tilde{\beta},\tilde{\theta}=\operatorname*{arg\,min}_{t,\beta,\theta}\sum_{i=1}^{N}w_{i}(\|N_{i}-\hat{N}_{i}\|_{2}+\|S_{i}-\hat{S}_{i}\|_{2}),$ (7)
where $w_{i}$ denotes the confidence of $i$-th view, $\hat{N}_{i}$ and $\hat{S}_{i}$ represent the SMPL-X normal and silhouette renderings from predefined views.
Remeshing with differentiable rasterization. Given the initial human prior, we utilize differentiable rasterization to carve the details based on observational normal maps. While a common approach involves adding per-vertex displacement to the coarse canonical mesh, this method encounters difficulties when modeling complex details, such as loose clothing. To address this limitation, we directly optimize the SMPL topology, encompassing both vertex positions $V$ and face edges $F$. The optimization procedure iteratively applies vertex displacement and remeshing to the triangle mesh, utilizing the optimizer proposed in [30]. The optimization objective can be written as
$\tilde{V},\tilde{F}=\operatorname*{arg\,min}_{V,F}\sum_{i=1}^{N}w_{i}(\|N_{i}-\hat{N}_{i}\|_{2}+\|S_{i}-\hat{S}_{i}\|_{2})+\lambda\sum_{j}(n_{j}-n_{j}^{\text{neig}}),$ (8)
where $w_{i}$ denotes the confidence of $i$-th view, $\hat{N}_{i}$ and $\hat{S}_{i}$ represent the normal and silhouette renderings from predefined views, $n_{j}$ and $n_{j}^{\text{neig}}$ denote the vertex normal and the average normal of neighboring vertices. The regularization weight $\lambda$ is set to $0.02$. We execute $700$ optimization steps to achieve optimal performance. Following the mesh optimization, we employ Poisson reconstruction [17] to complete minor invisible areas, such as the chin. Additionally, following [43], we offer the option to replace the hands with the estimated SMPL-X results to enhance visual quality.
Appearance fusion. Upon obtaining the 3D geometry, our objective is to derive the high-fidelity texture matching the reference image. Despite the availability of multiview images, direct projection onto the mesh results in conspicuous artifacts, arising from the cross-view inconsistency and inaccurate foreground segmentation. To overcome this, we perform texture fusion utilizing the aforementioned differentiable rendering. Specifically, we optimize the per-vertex color $VC$ by minimizing
$VC=\arg\min_{vc}\sum_{i=1}^{N}w_{i}\|C_{i}-\hat{C}_{i}\|_{2},$ (9)
where $C_{i}$ and $\hat{C}_{i}$ represent the rendered and generated color images, respectively. In the majority of cases, this color fusion pipeline suffices to generate high-quality appearances. However, certain areas may remain unobserved from the predefined six viewpoints. Thus, we finally compute a visibility mask and perform topology-aware interpolation based on KDTree, ensuring comprehensive texture coverage.
4 Experiments
   Input     PaMIR      PIFuHD         GTA         SiFU        HiLo        ECON          Ours
Figure 4: Geometry comparison of PSHuman with Implicit and Explicit methods for 3D human inference from in-the-wild images. Existing methods often struggle with complex poses and loose clothing, leading to issues such as absent body parts, disrupted clothing, and a lack of fine details. In contrast, PSHuman provides a complete shape, detailed facial features, and natural-looking clothing folds. Following [43], we substitute the hands with SMPL-X models to enhance visual quality.
Training and evaluation details. PSHuman builds upon the open-source pre-trained text-to-image generation model, SD2.1-unclip [34]. Our training is conducted on a cluster of 16 NVIDIA H800 GPUs, with a batch size of 64 for a total of $30,000$ iterations. We adopt an adaptive learning rate schedule, initializing the learning rate at 1e-4 and decreasing it to 5e-5 after $2,000$ steps. The entire training process spans approximately three days. Regarding the reconstruction module, we perform SMPL-X alignment, geometry optimization, and texture fusion for $700$, $100$ and $100$ steps, respectively, with corresponding learning rates of $0.3$, $0.001$, and $0.0005$. For appearance evaluation [51], we render color images from four viewpoints at azimuths of {$0^{\circ},90^{\circ},180^{\circ},270^{\circ}$} relative to the input view.
Dataset. We conduct experiments on widely used 3D human datasets, including THuman2.1 [47], CustomHumans [12] and CAPE [25]. Specifically, our training dataset comprises $2,385$ scans from THuman2.1 and $647$ scans from CustomHumans. These datasets are selected due to their provision of SMPL-X parameters. For quantitative evaluation, we utilize the remaining $60$ scans ($0447$-$0486$, $0492$-$0511$) from THuman2.1 and $150$ scans from CAPE. Following ICON’s partitioning criteria, we subdivide CAPE into ”CAPE-FP” (50 samples) and ”CAPE-NFP” (100 samples) to assess generalization in real-world scenarios.
Metric. To assess reconstruction capability, we employ three primary metrics: 1-directional point-to-surface (P2S), $L_{1}$ Chamfer Distance (CD), and Normal Consistency (NC). For appearance evaluation, we utilize peak signal-to-noise ratio (PSNR[41], structural similarity index (SSIM[48], and learned perceptual image patch similarity (LPIPS[49].
Table 1: Quantitative comparison of geometry quality. For the setting of ‘w/o SMPL-X body prior’, we utilize PIXIE to estimate SMPL parameters for other baseline methods while omitting SMPL estimation for our approach. Specifically, we retrain the diffusion model by removing the SMPL-X conditioning and initialize human mesh with a unit sphere during mesh carving. For ‘w/ SMPL-X body prior’, ground-truth SMPL-X models are used to avoid the impact of pose estimation errors on the evaluation. The units for Chamfer and P2S are in cm. The top two results are colored as first second.
CAPE-NFP CAPE-FP THuman2.1
Method Venue Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$
w/o SMPL-X body prior
PIFu ICCV 2019 3.2524 2.5469 0.7624 1.8367 1.7582 0.8573 1.2071 1.1299 0.7681
PIFuHD CVPR 2020 2.9749 2.3677 0.7658 1.5211 1.4834 0.8712 0.9935 0.9647 0.7890
PaMIR TPAMI 2021 7.1577 3.3832 0.6345 6.0114 3.2877 0.6737 1.0875 1.0144 0.7939
ICON CVPR 2022 2.6983 2.3911 0.7958 2.1331 2.0359 0.8364 1.1199 1.0925 0.7810
ECON CVPR 2023 3.1086 2.6044 0.7722 2.5394 2.4336 0.8128 1.2500 1.1469 0.7643
GTA NeurIPS 2023 2.7387 2.4722 0.7875 2.2543 2.1889 0.8247 1.0612 1.0389 0.7857
SIFU CVPR 2024 2.7884 2.4792 0.7877 2.1695 2.1107 0.8310 1.0774 1.0586 0.7871
HiLo CVPR 2024 2.6507 2.3037 0.7987 2.2735 2.1345 0.8308 1.1241 1.0519 0.7784
SITH CVPR 2024 2.8735 2.1226 0.7804 2.1140 1.6754 0.8337 0.9661 0.9034 0.7832
Ours - 2.1625 1.6675 0.8226 1.3615 1.1308 0.8844 0.6609 0.5993 0.8310
w/ SMPL-X body prior
ICON CVPR 2022 1.5511 1.1967 0.8572 0.9951 0.8864 0.9190 0.6146 0.5934 0.8493
ECON CVPR 2023 1.8524 1.5706 0.8392 1.1761 1.1352 0.8969 0.6725 0.6331 0.8362
GTA NeurIPS 2023 1.8853 1.4902 0.8260 1.1484 0.9914 0.9011 0.5791 0.5587 0.8491
SIFU CVPR 2024 1.5742 1.2777 0.8529 1.0535 0.9674 0.9024 0.5754 0.5576 0.8500
HiLo CVPR 2024 1.5613 1.2146 0.8547 1.1246 0.9847 0.9031 0.5977 0.5892 0.8405
SITH CVPR 2024 1.8118 1.5201 0.8345 1.1839 1.1573 0.8870 0.6474 0.5810 0.8264
Ours - 0.9688 0.8675 0.8799 0.7811 0.6984 0.9136 0.4399 0.4077 0.8504
4.1 Comparisons
Baselines. We conducted a comprehensive comparison of our method against state-of-the-art single-view human reconstruction approaches, including PIFu [35], PIFuHD [36], PaMIR [54], ICON [42], ECON [43], GTA [50], SiFU [51], HiLo [45], and SiTH [13]. For SMPL-based methods, we utilize PIXIE [46] for estimation. We also report the results with ground-truth SMPL-X to isolate the impact of pose estimation errors.
Comparison of geometry quality. Leveraging consistent multiview images, our method exhibits superior geometric quality compared to existing approaches, particularly in scenarios without SMPL-X body prior (Tab. 1). Unlike other template-based methods, which are susceptible to SMPL-X prediction errors, our method supports template-free training, thereby offering enhanced generalization capability. When incorporating the body prior, our method consistently outperforms previous works, demonstrating unprecedented accuracy on complex posed humans. The qualitative comparison in Fig. 4 also showcases the superiority of PSHuman, featuring with complete shape, detailed face and natural-looking clothing folds.
Table 2: Quantitative comparisons of face reconstruction.
Method Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ECON 0.624 0.570 0.837 - - -
SIFU 0.535 0.527 0.853 18.86 0.790 0.093
SITH 0.610 0.563 0.858 17.93 0.827 0.110
w/o local 0.524 0.503 0.867 19.67 0.832 0.093
w/o noise blender 0.447 0.422 0.904 20.85 0.877 0.075
Ours 0.423 0.397 0.924 20.97 0.896 0.071
Table 3: Quantitative comparison of appearance rendering on THuman2.1 subset.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
PIFu 19.3957 0.8327 0.1001
PaMIR 19.4130 0.8324 0.0988
GTA 19.6071 0.8338 0.0989
SIFU 19.4417 0.8307 0.0985
SITH 18.4580 0.8200 0.1004
Ours 20.8548 0.8636 0.0764
Table 4: Evaluation of robustness to SMPL-X estimation on THuman2.1 subset.
Method Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$
ICON 0.7827 0.6463 0.8401
ECON 0.8022 0.6742 0.8327
GTA 0.6631 0.6473 0.8368
SIFU 0.6672 0.6488 0.8302
SITH 0.6427 0.6393 0.8241
Ours 0.5574 0.5377 0.8417
Comparison of appearance quality. Quantitative evaluations in Tab. 4 reveal that PSHuman outperforms existing methods across multiple metrics, achieving the highest PSNR, SSIM as well as the lowest LPIPS. Qualitatively, as illustrated in Fig. 5, PSHuman produces highly consistent appearances on novel viewpoints, including natural and realistic reconstruction for posterior regions. In contrast, existing methods exhibit various limitations such as blurred colors and inconsistent artifacts in unseen views.
Comparisons of face quality. To highlight the effectiveness of our introduced cross-scale diffusion for face reconstruction, we use the head vertices of SMPL-X to crop the reconstructed head following ECON. Specifically, we first construct a KD-tree based on SMPL-X to query the generated mesh, subsequently filtering out the vertices adjacent to the head of SMPL-X. Tab. 2 presents the quantitative comparisons with SOTA methods.
  Input   PaMIR      GTA        SIFU  [^2] We only compare with the coarse texture because the refinement code is unavailable.       SiTH     Ours
Figure 5: Appearance comparisons with methods which produce texture. Our method could reconstruct realistic and reasonable appearance of side and back views.
4.2 Ablation Study
Effectiveness of SMPL-X condition. In Fig. 1, we show the geometry reconstructed by the models trained without SMPL-X condition and with SMPL-X condition. In Fig. 1(a), it is observed that the naive diffusion model struggles to ‘imagine’ the pose of a bending human image. Conversely, the SMPL-X provides a strong pose prior to guide the model to generate reasonable side views, leading to better reconstruction. In Fig. 1(b), the diffusion model fails to generate consistent multiple views due to self-occlusion, resulting in artifacts near the arm regions. The SMPL-X guidance effectively enhances consistency, facilitating the complete human body.
      Input      w/o local   w/o noise blendering   Ours
Figure 6: Ablation study of the cross-scale diffusion (CSD). The CSD allows sharp face recovery and keeps the identity consistent with the reference input.
Effectiveness of cross-scale diffusion (CSD). In Tab. 2, we provide the results by removing the local face branch (w/o local) and noise blending (w/o noise blending), respectively. Our method, incorporating both components, achieves the highest performance, as shown in Fig. 6. Notably, the setting without noise blending also generates the local face image. However, the reconstructions exhibit minor artifacts or over-smoothness. We attribute it to the inconsistency among global and local images. In contrast, the noise blending allows the information exchange, thereby enhancing overall consistency.
    Input      NeuS     BiNI     SMPLX-D  Naive Remesh  Ours
Figure 7: Ablation of our human carving module.
Figure 8: Visualization of mesh carving of a posed human image.
Effectiveness of mesh carving module. We assess the efficacy of our reconstruction module by substituting the remeshing step with alternative methods, specifically NeuS and BiNI. As illustrated in Fig. 7, the resulting geometries exhibit notable deficiencies or failures to capture fine geometric details. Note that we employ the normal maps, generated by our diffusion model, across all methods to mitigate potential errors arising from normal prediction discrepancies. Moreover, “naive remeshing” refers to remeshing with SMPL-X initialization but without multiview-guided SMPL-X alignment, resulting in subtle artifacts caused by misalignment between the initial SMPL-X mesh and the multi-view observations. Our reconstruction module effectively addresses these issues. Finally, we show an example across remeshing process for better understanding in Fig. 8.
Robustness to SMPL-X estimation. We assess the robustness of template-based approaches to SMPL-X estimation errors in Tab. 4. Following SIFU, we introduce random noise with a variance of $0.05$ to both the pose and shape parameters of the ground-truth SMPL-X model. The results demonstrate the robust reconstruction capabilities of our approach. Furthermore, the efficacy of our method in real-world scenarios is evidenced by the additional results presented in supplementary materials.
Comprehensive quantitative ablation. We further conducted comprehensive ablation studies on a subset of $20$ samples from the “CAPE-NFP” dataset. Tab. 5 quantitatively illustrates the impact on Chamfer Distance when individual components are removed or replaced. It is observed that the SMPL-X condition contributes significantly to reconstruction accuracy. While CSD yields a modest reduction in geometric error, it substantially improves visualization quality and identity fidelity, as evidenced in Fig. 6. Furthermore, our reconstruction method, which employs SMPLX-guided differentiable remeshing, demonstrates superior reconstruction performance compared to the BiNI and inpainting pipeline utilized in ECON.
Table 5: Comprehensive ablation study of core designs w.r.t full body reconstruction performance.
[…]
CSD SMPLX-Cond. Remeshing SMPLX-ECON SMPLX-Remeshing
1.4920
1.4370
1.0938
1.2630
0.9597 (Ours)
5 Conclusion
A. Inaccurate pose B. Stitching artifacts C. Wrong generation    D. Loose hair
Figure 9: Failure cases of PSHuman.
Limitations. Although PSHuman achieves high-quality single-view human reconstruction, it shares certain limitations with previous template-based approaches as shown in Fig. 9. First, pose estimation errors (A, B) have a cascading effect on subsequent generation and reconstruction, impacting overall accuracy. In addition, wrong novel-view generation (C) may result in unreasonable geometry. Finally, diffusion models struggle to generate consistent subtle details, such as loose hair and hands (D), which results in suboptimal reconstruction.
Conclusion. We present PSHuman, a novel framework that significantly improves geometric and appearance quality in single-image human reconstruction. We investigate direct multiview human generation conditioned on SMPL-X, enabling explicit and robust human reconstruction. Our body-face cross-scale diffusion model enhances the modeling of high-fidelity 3D human faces, while our multiview-guided explicit carving module ensures intricate details from generated images. Experiments demonstrate that PSHuman’s superiority against existing methods.
Acknowledgement. The research was supported by Theme-based Research Scheme (T45-205/21-N) from Hong Kong RGC, Generative AI Research, Development Centre from InnoHK, NSFC (No.62171255) and Guoqiang Institute of Tsinghua University (No.2021GQG0001).
Supplementary Material
6 Discussions about face-body cross-scale diffusion
Difficulty in implementing dependent forward process. In the dependent forward process $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$, we know that the face region of $x^{B}$ corresponds to $x^{F}$. Since we have defined $p(x^{F}_{t}|x_{t-1}^{F})$ by adding noises to $x^{F}_{t-1}$, it is natural to get $x^{B}_{t}$ by replacing the pixel values in the face region of $x^{B}_{t}$ with $x^{F}_{t}$ and just adding noises to the remaining image regions of $x^{B}_{t-1}$. However, since we adopt a latent diffusion model (Stable Diffusion) Rombach et al. [33] here, the pixels of tensors in the latent spaces are not independent of each other so the replacing operation is not valid here. This brings difficulty in separating the face regions in the latent space to explicitly implement the dependent forward process for adding noises.
Rationale of approximated forward process. Our rationale for adding noises to the face and the body separately is that the process is similar to multiview diffusion. We can regard the face image and the body image as just two images captured by cameras with different camera positions and focal lengths. In this case, the body-face cross-scale diffusion is a special case of multiview diffusion. In a multiview diffusion, we add noises to multiview images separately so that we can also add noises to the body image and face image separately but consider the dependence in the reverse process.
7 Implementation Details
Preprocessing. Our training datasets include scans from THuman2.1 and CustomHumans. For each human model, and the corresponding SMPL-X model, we render 8 color and normal images with alpha channel around the yaw axis, with a $45^{\circ}$ interval and a resolution of $768\times 768$. Due to the random face-forward direction, we employ insightface Deng et al. [5] for face detection, utilizing only viewpoints containing clear facial characteristics for training.
Choice of generated views. As mentioned in the main paper, PSHuman generates 6 color and normal images from front, front-right, right, back, left, and front-left views for the trade-off between effectiveness and training workload. To guarantee the generation alignment, we horizontally flip the left and back views during training. In Fig. 11, we present the results reconstructed using only two-view (front and back) or four-view (front, right, back, left) normal maps. Since there is a lack of depth in information, optimizing geometry with fewer views leads to severe artifacts, such as incomplete or unnatural human structures. In contrast, it is evident that the artifacts are reduced when using six views.
Diffusion block. As illustrated in Fig.3(b) of the main paper, our diffusion block comprises two branches. The local diffusion inherits from stable diffusion (SD2.1-Unclip) [34], including self attention, cross attention and feed-forward layers, while the global attention contains an additional multi-view attention layer introduced in Era3D [21]. The global attention is conditioned on the local branch via the noise blending layer. We feed the embeddings of text prompt ”a rendering image of 3D human, [V] view, [M] map.” into the denoising blocks via cross attention, where [V] is chosen from ”front”, ”front right”, ”right”, ”back”, ”left”, ”front left”, ”face” and [M] represents ”normal” or ”color”.
Inference details. Given a human image, we first remove the background with rembg [7] and then resize the foreground to 720$\times$720. Finally, we pad it to 768$\times$768 and set the background to white. Due to the alignment between of processed input image and the generated front color image, we use the former and other generated images in the following reconstruction.
8 More experiments
Inference time. In Tab. 6, we report the detailed inference time of the whole pipeline, including preprocessing (SMPL-X estimation and SMPL-X image rendering), diffusion, geometry reconstruction (SMPL-X initialization and remeshing) and appearance fusion.
Table 6: Inference time of the reconstruction module.
Pipeline Pre-processing Diffusion Geo. Recon. Appearance Fusion
Time / s 7.2 17.6 23.3 6.0
Table 7: User study w.r.t reconstruction quality and novel-view consistency.
Method PIFuHD PaMIR ECON GTA SiTH Ours
Geometry Quality 1.55 1.96 3.72 2.11 2.72 4.71
Appearance Quality - 1.42 - 2.65 2.82 4.59
Geometry Consistency 1.69 1.76 2.48 2.33 2.79 4.61
Appearance Consistency - 1.77 - 2.16 2.73 4.68
Figure 10: Qualitative comparison with optimization-based methods. We demonstrate the results of (a) Magic123, (b) Dreamgaussian, (c) Chupa, (d) TeCH and (e) Ours.
User study Given the limitations of quantitative metrics in assessing the realism and consistency of side and back views reconstructed from single-view input, we conducted a comprehensive user study to evaluate the geometry and appearance quality of five SOTA methods. Specifically, we collect 20 in-the-wild samples and 20 cases from SHHQ fashion dataset for evaluation. Following HumanNorm [14], we invite 20 volunteers to evaluate the color and normal video rendered from the reconstructed 3D humans. Participants were instructed to score each model on a 5-point scale (1 being the worst and 5 being the best) across four key dimensions:
  • To what extent does the human model exhibit the best geometry quality?
  • To what extent does the human model exhibit the best appearance quality?
  • To what extent does the novel view’s geometry of the human body align with the reference image?
  • To what extent does the novel view’s appearance of the human body align with the reference image?
For methods that do not produce texture (PIFuHD and ECON), we only compare the geometry quality and consistency. The results in Tab. 7 indicate that our method represents a significant advancement against SOTA methods, offering superior performance in both geometry and appearance reconstruction, as well as consistency across novel viewpoints.
Comparison with optimization-based methods. To assess the efficacy of our approach relative to optimization-based methods, we conducted a comparative analysis of PSHuman against several SDS-based techniques, Magic123, Dreamgaussian, Chupa, and TeCH. Following SiTH, we adopt the pose and text prompt generated by [20] as condition inputs due to the lack of direct image input support in Chupa. As illustrated in Fig. 10, Magic123 and Dreamgaussian exhibit significant limitations, primarily manifesting as incomplete human body reconstructions and implausible free-view textures. The reliance on text descriptions for conditioning proves insufficient for fine-grained control, resulting in geometries that deviate substantially from the reference inputs. TeCH, a method specifically designed for human reconstruction from a single image, while capable of producing complete human shapes, struggles with severe noise in geometric details and over-saturated textures. These artifacts are characteristic challenges inherent to SDS-based methodologies. In contrast, PSHuman demonstrates superior performance by directly fusing multi-view 2D images in 3D space, enabling the preservation of geometry details at the pixel level while circumventing unrealistic texture. Note that TeCH requires $\sim$6 hours for optimization, PSHuman generates high-quality textured meshes within merely 1 minute. We refer readers to Fig. 19 and Fig. 20 for more results generated by PSHuman.
    Input     view=2       view=4       view=6 (Ours)
Figure 11: Ablation of view number. Since normal maps lack depth information, optimizing geometry by only two or four views leads to an incomplete or unnatural human structure.
Capability of handling occlusion. We present the generated normal maps (back, left, and right views) and corresponding meshes of in-the-wild samples with various self-occlusion, as demonstrated in Fig. 17. To further illustrate the robustness of our approach, we also include examples of object-occluded scenarios in Fig. 12. The results show that our diffusion model can infer the correct human structure under both self-occlusions and object occlusions, enabling the reconstruction of high-quality 3D meshes even under such challenging conditions.
Figure 12: Reconstruction quality on object-occluded images.
Robustness to SMPL-X estimation. The SMPL-X serves as a coarse anatomy guide, only required to be reasonably overlayed with the human body. Thus, our method could handle estimation error (Fig. 13) to some extent and generalize to children or the elder in Fig. 14.
Figure 13: Robustness to SMPL-X estimation errors.
Figure 14: Performance with out-of-distribution pose estimation, like children and the elder.
Robustness to lighting. By incorporating varying lighting conditions using HDR maps from Poly Haven during training, our model demonstrates robustness to lighting variations, as illustrated in Fig. 15.
Figure 15: Robustness to shading and strong light.
Comparisons of face normal estimation. As shown in Fig. 16, our local face diffusion model generates facial normal images with significantly enhanced fine-grained details compared to ECON [43] and SAPEIN-2B [18].
     Input         Ours     Sapiens-2B       ECON
Figure 16: Comparisons of face normal estimation.
Generalization on anime characters. Our model, trained with only realistic human scans, exhibits excellent generalization on anime or hand-drawn style character images, as shown in Fig. 18. This is because our method is adapted from the Stable Diffusion [34] model, which has been trained on images of various styles. Thus, our method maintains the ability to generalize images of different domains.
9 Ethics statement
While PSHuman aims to provide users with an advanced tool for single-image full-body 3D human model reconstruction, we acknowledge the potential for misuse, particularly in creating deceptive content. This ethical concern extends beyond our specific method to the broader field of generative modeling. As researchers and developers in 3D reconstruction and generative AI, we have a responsibility to continually address these ethical implications. We encourage ongoing dialogue and the development of safeguards to mitigate potential harm while advancing the technology responsibly. Users of PSHuman and similar tools should be aware of these ethical considerations and use the technology in accordance with applicable laws and ethical guidelines.
Figure 17: Reconstruction quality on self-occluded images. We present the generated back, left, and right views of normal maps and corresponding meshes.
Figure 18: Generalization on anime characters. We present the generated multiview color and normal images and corresponding meshes (in blue).
Figure 19: More results on SHHQ dataset.
Figure 20: More results on in-the-wild data.